### v1

# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

Abstract.
The emergence of digital payment systems has transformed how individuals conduct financial transactions, offering convenience, security, and efficiency. One groundbreaking innovation making waves in the Indian financial landscape is the Unified Payments Interface (UPI), developed by the National Payments Corporation of India (NPCI). Existing work has explored how digital payments benefit a country’s economy and GDP. However, our study explores how the introduction of UPI has influenced spending behavior among Indian users on an ”individual” level. We gathered 235 valid survey responses encompassing diverse demographics and conducted semi-structured interviews with 20 survey respondents. Approximately 75% of the survey respondents reported increased spending due to UPI, with only 7% indicating reduced spending. Significantly, 91.5% of the respondents reported satisfaction with their UPI usage. Also 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management.
Digital Payments, Cashless Transactions, Unified Payments Interface (UPI), User Study
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer-Supported Cooperative Work and Social Computing; November 09–13, 2024; San Jose, Costa Ricabooktitle: CSCW’24: ACM Conference on Computer-Supported Cooperative Work and Social Computing, November 09–13, 2024, San Jose, Costa Ricaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
The term ”E-payment” is comprehensive, illustrating diverse proportions of E-payments through different channels. Despite attempts to define electronic payment, a clear, universally accepted definition is still lacking. An E-payment is generally understood as an economic situation where the buying and selling of goods are handled without physical cash, often facilitated through E-payment or checks (Pau, 2012). Digital payments have become integral to modern financial landscapes, offering swift, secure, and convenient electronic transactions that transcend traditional methods. This global shift towards digital transactions has been accelerated by the widespread adoption of internet and mobile technologies, reshaping financial practices across the world (Alkhowaiter, 2020). In 2022 alone, the volume of real-time digital transactions reached an unprecedented 195 million (89., 2023). Notably, 64 % of adults globally engaged in at least one digital payment transaction, with high-income economies leading at 95% (glo, 2021).
India, in particular, has emerged as a global leader in digital transactions, accounting for 46 % of the total global digital transactions in 2022 (89., 2023). At the forefront of this digital revolution is the Unified Payments Interface (UPI), launched in 2016 by the National Payments Corporation of India (NPCI). UPI has redefined financial transactions by enabling instant, real-time inter-bank transfers (Mistry, 2023). As of August 2023, UPI boasts the participation of 484 banks and has facilitated a staggering 10 billion transactions, totaling Rs 18,22,949.42 Crore(189 billion USD) (npc, 2023a). In India, 84 % of all digital transactions in 2022, amounting to 1.79 trillion dollars, were channeled through UPI (npc, 2023b). UPI’s rapid proliferation and high penetration in the Indian payment ecosystem underscore its transformative impact (Mistry, 2023).
Despite UPI’s monumental success, a critical research gap exists regarding its influence on users’ spending habits. Our study addresses this crucial gap in existing research on UPI, specifically focusing on the impact of UPI on users’ spending habits. While previous studies and works (Sreenu, 2020) acknowledge UPI’s success in reshaping digital payments, they fall short of providing detailed insights into how this technology influences ”individual” spending behaviors. The existing literature often emphasizes macro-level implications and adoption rates. Multiple questions about changes in expenditure preferences, budgeting strategies, investment choices, long-term financial goals and overall financial behaviors remain unexplored, hindering a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors. It adds valuable knowledge to the broader discourse on digital payment systems, stressing the importance of considering individual-level implications for a more holistic comprehension of financial technology’s transformative effects. Moreover, there’s a notable absence of research delving into potential variations in UPI’s impact across different demographic groups. Age, profession, and socio-economic status are overlooked in understanding how individuals interact with UPI and manage their finances.
We also examine users’ perceptions of security, investigating how trust in UPI’s transactional security affects their expenditure patterns. Additionally, the role of user trust in technology, particularly in UPI adoption, remains underexplored, prompting our study to unravel the intricate dynamics between trust, technology adoption, and financial decision-making. Recognizing the transformative nature of UPI, we also investigate the relationship between users’ financial literacy levels and their utilization of UPI, shedding light on how understanding financial principles shapes spending choices. Furthermore, our research extends its focus to entrepreneurs and businesses, probing into the effects of UPI adoption on their financial practices and business models. Another unique contribution of our research is analyzing UPI transactions’ integration with budgeting applications, employing a competitive lens to understand its impact on users’ awareness of spending patterns and financial discipline. Lastly, our study explores the dynamic nature of user satisfaction with UPI, considering system updates, security enhancements, and evolving user experiences. Collectively, our research endeavors aim to understand UPI’s influence on diverse aspects of users’ financial behaviors, contributing valuable insights to the evolving landscape of digital payments in India.
The significance of this study extends beyond the realm of finance. In Human-Computer Interaction (HCI), understanding how users interact with financial technologies is essential for designing user-friendly and effective interfaces. This study contributes valuable insights to HCI researchers, financial institutions, and policymakers, guiding the development of digital payment systems that align with users’ financial needs and behaviors.
In addressing the above-mentioned gaps, this research employs a mixed-methods approach, encompassing qualitative and quantitative assessments. We gathered 276 responses covering diverse demographics and backgrounds using a survey form, out of which 235 were valid responses. Additionally, we also conducted semi-structured interviews with 20 survey respondents. This study aims to address the following research questions (RQs):
  1. (1)
    RQ1: How does the introduction of UPI influence Indian users’ spending habits, and what are the observed changes: increased, decreased, or unchanged spending patterns?
  2. (2)
    RQ2: How do users perceive the convenience and efficiency of UPI transactions compared to traditional cash transactions, and to what extent do these perceptions shape their spending behavior?
  3. (3)
    RQ3: To what extent does UPI adoption vary across demographic groups, such as age and profession, and how do these variations influence users’ financial practices, including spending habits and long term financial goals?
Among the participants we surveyed, 74.2% of the participants confirmed increased spending due to UPI, with only 7% confirming reduced spending. Significantly, 91.5% of respondents reported satisfaction with their UPI usage. Also 95.2% of the users found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. To the best of our knowledge, this is the first user study of UPI’s impact on individual spending behaviour in India.
2. Related Work
In recent times, there has been a growing interest in previously overlooked countries such as the BRIC nations, including Brazil, Russia, India, and China, where rapid economic and technological development is occurring (Kumar et al., 2011). Mobile payment systems are progressively influencing people’s socio-economic lives in South Asia (Asongu and Boateng, 2018; Asongu and Odhiambo, 2019; Ferreira and Perry, 2019; He et al., 2023; Lewis and Perry, 2019; Vashistha et al., 2019). India topped the list of the five countries, namely Brazil, China, Thailand, and South Korea, for digital payments by the end of 2022 with 89.5 million transactions (89., 2023). The sheer nature of mobile payments and their ubiquity, along with convenience, helps include individuals, eventually benefiting the country’s economy. Branchless banking (mobile money) apps are responsible for almost 30% of the GDP in various countries (ken, 2013; Ferreira and Perry, 2019).
Historically, India has been characterized by high cash usage, with a substantial portion of the workforce engaged in the informal sector relying on cash transactions (Ghatak, 2017; Pal et al., 2018a). However, the government’s ”Digital India” vision, coupled with the rapid expansion of internet access, has propelled the adoption of digital payments. This transformation, from a cash-driven economy to a significant shift towards digital payments is also driven by various significant events like Demonetization and the COVID-19 pandemic. The rise of digital wallets and UPI platforms, including BHIM UPI (bhi, 2016), Paytm (pay, 2010), PhonePe (pho, 2015), and Google Pay (goo, 2012), etc. have played a pivotal role in reshaping the country’s financial landscape. Internet penetration in India increased from 14% in 2014 to over 46% in 2021, primarily attributed to affordable data tariffs introduced by regional mobile operators in 2016, with prices as low as Rs 15 (0.21 USD) per gigabyte (GB) of mobile data. As a result, India ascended to the second-largest global user base of the internet by 2018 (2ND, 2019; Jio, 2019).
Key events, such as Demonetization in 2016, in which, according to the government’s decision, most of the nation’s banknotes were declared out of legal tender aimed to combat corruption and terrorism (ter, 2017; Pal et al., 2018b; Kameswaran and Hulikal Muralidhar, 2019), further accelerated the shift towards digital payments. The COVID-19 pandemic acted as another catalyst, prompting the Reserve Bank of India to advocate for reduced cash usage to minimize physical contact, significantly increasing digital transactions. According to the World Bank, over 80 million Indian adults made their first digital merchant payment after the pandemic’s onset, further solidifying India’s position as a global digital payment leader (wor, 2022). Studies have indicated the positive impact of electronic transactions on trade and economic development in India (Abr, 2004; muk, 2011). The benefits include transparency, accountability, and reduced cash-related fraud, contributing to overall economic growth (Shiva, 2017).
Despite progress, challenges remain, as only 35% of Indian adults use bank accounts for digital payments (glo, 2021). The low percentage (approx 5% of the Indian population) of income tax contributors emphasizes the potential role of digital payments in increasing tax collection, as highlighted by previous research (bax, 2005). The extensive penetration of smartphones in India, with 1.2 billion active mobile phone users, underscores the technological underpinnings of this digital transition (MIN, 2022). In 2016, a study by Goel (goe, 2016) found that technological advancements have played a significant role in empowering and introducing innovations in E-payment systems. Promoting digital payments is also associated with reducing cash-related crimes, as individuals tend to hold less physical cash, thus discouraging robbers from pursuing such crimes.
Research in the field has extensively explored the impact of mobile payment systems on economies and societies, focusing on their transformative effects. Notably, people from different countries and different strata of society manage finances differently (Fran Bennett, 2009; Kaye et al., 2014; Vines et al., 2011). This motivated us to understand the possible changes in their spending due to UPI and how Indian users keep track of their finances. Over the past decade, these systems have helped raise the standard of living and have revolutionized how money is used in developing economies.
While previous research has broadly addressed the impact of mobile payment systems, our study focuses specifically on the financial behaviors of Indian users, particularly those using UPI. While the existing literature has focused on macro-level implications, multiple questions about changes in expenditure preferences, budgeting strategies, long-term financial goals and overall financial behaviors remain unexplored. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors, thus, giving a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. Additionally, our research pays particular attention to the concerns and preferences of users, shedding light on factors that may influence the widespread acceptance of mobile payment systems in the region. Moreover, our work also aims to understand user trust in UPI technology. The study aims to uncover the factors that encourage users to trust the UPI technology, considering the backdrop of rising scams (sca, 2021, 2023). By delving into this aspect, we aim to provide insights beyond transaction numbers, contributing to a more comprehensive understanding of India’s evolving digital payments landscape.
Evidence suggests that individuals often under-save, fail to invest wisely, and are frequently indebted (stu, 2012). Increasing financial literacy improves financial decision-making, leading governments worldwide to recognize it as a critical life skill, launching financial education initiatives to help young people acquire this skill (asi, 2014; rbi, 2020). Multiple studies have highlighted the importance of financial inclusion and literacy, with one study pointing out how roadblocks to digital or financial literacy impede digital financial inclusion (Srivastava, 2022). A study by Agarwalla et al. (iim, 2014) focused on Indian working youth, stating that focused intervention is needed to improve the ”financial attitude” of men and the ”financial behavior” of women. French et al. (French et al., 2021) examined the effectiveness of personal finance smartphone apps in enhancing financial capability for low-income individuals by developing four specific apps. While cashless transactions promote technology literacy (tan, 2019), there is a visible gap in understanding how or if it promotes financial literacy. A financial activity study was carried out in 2014, tracking people’s transactional activity over time and interviewing them about their practices, understandings, needs, concerns, and expectations of present and future financial technologies to explore how people might interact with such digital payment systems (Lewis and Perry, 2019). As demonstrated by Lusardi et al. (llm, 2012), financial literacy training at an early stage in the working career significantly impacts financial outcomes due to compounded returns over a more extended period. Several studies worldwide indicate that a significant fraction of individuals have a low level of financial knowledge, associated with sub-optimal financial outcomes in various areas such as retirement planning (Lusardi and Mitchell, 2014; Lus, 2009; roo, 2011). In contrast to the existing studies, our study aims to explore how financial literacy impacts spending habits, especially with the onset of cashless transactions (UPI) in the context of India.
Finally, in this study, we conduct a comparative analysis of budgeting apps in India, an in-depth analysis of customer expectations from financial management or budgeting apps, and an understanding of specific user preferences that can act as a baseline for concerned developers, authorities, and stakeholders.
3. Methodology
3.1. RESEARCH DESIGN - MIXED DESIGN
In our research design, we employed a mixed methods (mul, [n. d.]) approach to comprehensively investigate the impact of India’s Unified Payments Interface (UPI) on individual spending behaviors and financial practices. This approach allowed us to gain a holistic understanding of how UPI has influenced users’ spending habits. We utilized two primary data collection techniques: surveys and interviews.
We employed a random sampling method for the survey component, targeting individuals aged 18 and above who may/may not have used UPI for payments. Our survey was conducted through Google Forms to ensure easy accessibility and respondent anonymity (by not asking for their personal information like name, phone number, etc as seen in Section 7). It garnered 276 responses from a diverse participant pool spanning various age groups and occupations, including working professionals, students, and business owners. It was floated across various social media platforms and groups and no questions were marked as mandatory, however, respondents were encouraged to provide honest and accurate responses as much as possible.
To ensure the quality and relevance of the responses, we implemented stringent criteria for validity. Frivolous responses containing irrelevant information were excluded, as were responses that deviated from the expected format. For instance, in the question regarding the initiation of UPI usage (”When did you start using UPI?”), any response indicating usage before 2016 was deemed invalid, considering that UPI was publicly launched that year. This validation process resulted in 235 valid responses for inclusion in our research work. Furthermore, to maintain consistency, we addressed ambiguous responses to the question about the commencement of UPI usage. For instance, answers like ”2 years back” were standardized to 2021, aligning with the survey release date of 2023. By adopting these rigorous validation measures, we ensured a reliable and high-quality dataset for our analysis.
In addition to surveys, we conducted in-depth interviews to dive deeper into the experiences and perspectives of UPI users. We randomly selected interviewees from our survey respondents who expressed an interest in participating in discussions. We conducted 20 (13 Male/ 7 Female) interviews over Google Meet and the transcripts of these interviews were further analysed. The interviewees provided verbal consent for the interview.
3.2. ETHICAL CONSIDERATION
In conducting this research, we implemented various ethical considerations to uphold transparency and safeguard the privacy and well-being of participants. Our study’s research materials and protocols underwent thorough review and approval by our university’s Institutional Review Board (IRB). Before participating in interviews, all candidates received information about the study’s purpose, the voluntary nature of their involvement, and the assurance of anonymity and confidentiality. We also acknowledge the significance of capturing diverse perspectives and experiences of individuals using UPI and its influence on their spending behavior. To ensure inclusivity, we actively sought participants from various backgrounds and genders. Verbal consent was obtained from participants for recording the interviews.
3.3. LIMITATION
One notable constraint of this study is the geographic scope of our data collection efforts. The research exclusively focused on respondents from urban areas in India, where technology adoption and usage, including electronic payment methods like UPI, tend to be higher and well integrated into daily life. Urban living typically involves residing in densely populated areas with access to advanced infrastructure and services, while rural living often means residing in less densely populated areas with a focus on agriculture and simpler infrastructure. As a result, the findings may not encompass the diverse spending habits of rural areas. They may not accurately represent the impact of UPI on the spending habits of living in rural areas in India.
4. Evaluations
4.1. QUANTITATIVE ANALYSIS ( ANALYSIS OF SURVEY RESULTS)
Our survey received 235 valid responses from a diverse range of respondents from various demographics and backgrounds. Among the respondents, 51.6% identified as students, 42% as working professionals, and 4.7% as business owners. Regarding gender distribution, the data shows that 77% of respondents were male, 19.7% were female, and 3.3% preferred not to disclose their gender. Regarding the adoption timeline of UPI as shown in figure 1(a), a notable 12% of respondents started using it in 2016 after the ’Demonetization’ of Rs 500 (6.03 USD) and Rs 1000 (12.07 USD) notes to combat black money and counterfeiting. Followed by 7.60% in 2017, 13.60% in 2018, 15.60% in 2019, 16% in 2020, 17.6% in 2021, 11.20% in 2022, and 2.40% in 2023. This diverse range of adoption dates reflects users’ varying entry points into the UPI ecosystem, often aligning with significant events like demonetization and pandemic-related lockdowns.
To understand the impact of UPI adoption, 74.2% of the survey respondents admitted that UPI had led to increased spending, while only 7% said that it had decreased spending. Additionally, 18.8% of the survey respondents felt no changes in their spending habits after adopting UPI. Regarding satisfaction levels, as shown in figure 2(a), 55% of the survey respondents indicated they were highly satisfied with their overall UPI experience. Additionally, 36.5% reported satisfaction, and 7.7% remained neutral in their assessment. Therefore we can conclude that 91.5% (55% + 36.5%) of the respondents are satisfied with their overall experience using UPI for payments. When asked to rate their convenience with UPI for making payments on a scale from 1 (lowest) to 5 (highest), as shown in figure 4, the survey results indicate that 69.50% of respondents rated ”5” (Highly convenient), 25.7% rated ”4” (convenient), and 3.7% were ”3” (Neutral). Therefore we can conclude that 95.2% (69.5% + 25.7%) of the respondents find using UPI for payments convenient. Notably, only 2-3 respondents reported dissatisfaction with UPI as a payment method, indicating that most were satisfied with their experience with UPI usage.
Regarding daily spending using UPI, respondents reported a wide range, with amounts varying from Rs 50 (0.60 USD) to 1 lakh (1,202.66 USD). On average, 15.7% of respondents mentioned their daily average spending is around Rs 200 (2.41 USD). When respondents were asked about their frequency of using UPI for payments, as shown in figure 1(b), an overwhelming 81.1% indicated that they use it daily. Additionally, 15.6% reported using UPI weekly, while 1.1% said they use it monthly, 1.90% mentioned using it occasionally, and 0.30% mentioned rarely. This data underscores the widespread and frequent usage of UPI among the surveyed individuals.
When asked about exceeding their budget or overspending due to UPI usage, as shown in the figure 3(a), 59.8% of respondents acknowledged experiencing such situations, while 39.8% believed they had not. Additionally, respondents had diverse opinions when asked about the impact of using UPI on their budget adherence, as shown in figure 2(b); approximately 23% strongly agreed (rating it as 5), 23.3% agreed (rating it as 4), 26.7% remained neutral (rating it as 3), 15.6% disagreed (rating it as 2), and 11.5% reported strongly disagree (rating it as 1), indicating that UPI has had an impact on their ability to stick to their budget and control overspending.
Figure 1. (a) When did users start using UPI? (b) How often do users use UPI?
Figure 2. (a) Satisfaction Level with UPI (b) UPI Impact on Budget
Figure 3. (a) Budget exceedance due to UPI (b) How often do people keep track of transaction history?
Figure 4. Convenience of making payments using UPI
Respondents employed various methods to manage their expenses, including popular platforms like Splitwise (spl, 2011) and Paytm (pay, 2010), bank statements, and applications such as SBI YONO (yon, 2017) and Samsung Notes (sno, 2016). Other tools included Excel sheets and dedicated finance apps like MyMoneyy (mym, 2022), Honeybook (hon, 2013), Monefy (my, 2016), Payhawk (pay, 2018) and Money Manager (mm, 2020). Notably, many respondents mentioned that they need to check their messages or transaction history regularly. This observation sheds light on users’ varied financial management, literacy, and discipline approaches. Furthermore, participants provided additional comments, and some expressed their willingness to participate in interviews. These findings illustrate a wide range of responses and opinions among UPI users, with a significant proportion acknowledging increased spending and high overall satisfaction with the platform and its convenience.
4.2. QUALITATIVE ANALYSIS (ANALYSIS OF INTERVIEW RESULTS)
4.2.1. EXPLORING UPI USAGE FOR TRANSACTIONS: INSIGHTS AND PERSPECTIVES
Respondents shared valuable insights regarding their experiences with UPI usage in our research findings, revealing common themes, patterns, and occasional divergent viewpoints. Many respondents lauded UPI for its convenience as a payment method, emphasizing its ability to eliminate the need to carry money physically. This convenience streamlined small payments, rendering them more efficient. Moreover, the consistent praise for UPI transactions centered on their remarkable speed and efficiency. Respondents frequently described these transactions as exceedingly fast and convenient.
According to a State Bank of India report, for every Rs 1 (0.012 USD) increase in UPI, there is an 18 paisa (0.0022 USD) decline in debit card transactions, indicating that the number of ATM visits has decreased from 16 to 8 times a year (dho, 2023). Also, the fact the transaction creates a digital record helps them keep a record of transactions and financial documentation. These factors collectively contribute to the high satisfaction levels among UPI users
“I love UPI! It’s a fantastic way to spend money without dealing with change or carrying cash. You can make payments from just Rs 1 (0.012 USD) to a whopping 1 lakh rupees (12,028.79 USD) in one go. It’s one of the best things in this progressive era.”-[P1]
The cashless nature of UPI was another aspect that resonated with several participants, simplifying the payment process by obviating the requirement to carry tangible currency. Furthermore, certain respondents, particularly a lawyer and a clinic owner, emphasized how UPI had markedly enhanced their business operations by expediting payments. This improvement in efficiency contributed to a smoother and more streamlined financial process for these businesses.
“As a lawyer, frequent photocopies and official expenses often require change money. However, I typically only carry a little cash with me. This led to a time-consuming process, motivating me to embrace UPI for seamless small payments.”-[P2]
One salient feature of UPI is that it has no processing/transaction fees (for P2P transactions), unlike credit and debit cards. Also, processing fees may or may not be there, depending upon the transaction amount and other factors. This benefits not only the merchants as well as the customer.
”I encourage my customers to use UPI or cash for payments instead of credit or debit cards to avoid the 2-3% transaction fees, which either come from my earnings or are passed on to the customer, potentially harming my reputation. UPI offers a real-time, no transaction fees and transparent payment solution.”-[P3]
Many respondents believe that possessing financial proficiency is essential when utilizing UPI. Individuals must delve into personal finance education and exercise restraint in expenditures to mitigate these issues. Improving money management habits is imperative to maintain financial stability while utilizing UPI.
”Personally, there are occasions when I exceed my initial spending plans, but I rely on finance tracking applications to ensure I stay within my budget.”-[P4]
The adoption of UPI has not been without its share of risks and challenges.A recurring theme was the awareness of potential for scams and fraudulent activities and its rising numbers. Several respondents recounted incidents involving phishing attempts, where malicious actors impersonated their known ones or UPI platforms to deceive users into sharing sensitive information. Additionally, users recounted instances of phishing messages disguised as official UPI notifications, urging them to click on malicious links and disclose sensitive information such as OTPs (One-Time Passwords) and bank details. These fraudulent activities have raised questions about the vulnerability of users to cyber threats. To address these challenges, users can take proactive measures to protect themselves. Recommendations based on user experiences include avoiding sharing sensitive information, such as OTPs, with anyone, regularly updating UPI apps, enabling two-factor authentication, and being cautious when receiving unsolicited payment requests. According to the Ministry of Finance, more than 95,000 fraud cases linked to UPI transactions were recorded in 2022-23, an increase from 84,000 incidents in 2021-22 (sin, 2023).
“I once had a bad experience with UPI. I got a message asking me to send money, and they wanted my bank OTP. Sadly, I gave it to them, and things got messed up. They took a big chunk of cash, around 10,000 rupees, from my bank account. Fortunately, my father worked at a bank and helped fix it. I learned a harsh lesson about being careful with UPI and not sharing important info.”-[P6]
”Whether UPI is good depends on how careful you are with it and how well you handle your money. If you’re smart about it, it can be useful. For example, my friend helped me get my money back from a scammer using UPI, which was great. But you’ve got to be cautious and use it wisely, or it might not turn out well.”-[P5]
4.2.2. FACTORS INFLUENCING SPENDING BEHAVIOR
A few respondents strongly preferred using UPI as a payment option. It simplifies transactions, reduces dependency on cash, and encourages frequent expenditure. This shift contributes to a more efficient, cashless economy, revolutionizing how people handle their finances and everyday purchases.
”During and after COVID, more transactions moved online, making things even more convenient. This helped the economy grow. When I have less cash, I tend to spend less. But I spend more often with UPI because there’s no need to hold back. You don’t have to keep track of physical cash; you can check your bank balance and pay directly from your bank account.”-[P7]
“UPI has made transactions easier. Combined with the rise of Swiggy/Zomato/Zepto/Blinkit, etc., ordering online has become much more appealing and easy. Without UPI, I might not have been spending much on food online. It’s so easy; no need to think twice. Sometimes, I regret purchases because you only need a phone with wifi/mobile data”-[P8]
This digital payment method has reshaped how individuals perceive and manage their finances, sometimes leading to more frequent and impulsive spending patterns.
”Cash feels like something tangible, like giving away something valuable when we spend it. But with UPI, it’s just a number on the screen, and it doesn’t feel the same way.”-[P9]
“Running out of cash is never a problem now. Spending does increase as you do not feel the pinch of physical cash being spent. As you always carry your whole bank balance, spending increases.”-[P10]
“Digital money does not give you guilt when spending. When you spend paper money, you feel guilty.”-[P11]
Many people often struggle with regret over their financial choices, and the reasons behind this regret are complex. This regret often comes from spending too much, needing to be mindful of their expenses, and making less-than-ideal financial decisions. The issue’s core lies in the sharp difference between physical cash and the abstract nature of digital transactions (UPI payments). While cash represents a physical object that one can touch and watch disappear from their wallet, digital transactions involve only numbers on a screen, disconnected from the physical world. This lack of a tangible connection and the absence of a solid feeling about the money being spent lead to spending without much thought, creating a sense of emptiness later filled with regret and guilt.
When we asked users about their spending habits and where they generally use UPI, a diverse range of insights emerged. It was evident that UPI has found its way into various aspects of daily life, transforming how people manage their expenses. Users reported using UPI for essentials like rent, groceries, and bills, as well as for work-related expenses, leisure activities such as dining out and entertainment, education, online shopping, and even small everyday indulgences like street food and auto-rickshaw fares.
I have increased expenditure on minor items such as tea or lime soda! However, since most establishments in my area now support UPI payments, it’s convenient when I forget to carry cash. I seldom have any physical currency with me these days. -[P12]
Additionally, some users rely on UPI for critical business transactions, emphasizing its role in personal and professional financial activities. With the integration of UPI as a mode of payment in almost every app, UPI is ubiquitous and a definite game changer in the Indian economic landscape.
Most respondents agreed with the statement that using UPI affects their ability to stick to their budget. They either agreed or strongly agreed that UPI impacts their budgeting habits. This suggests that for many users, UPI does influence their spending behavior and financial discipline.
“UPI brought the revolutionary change in the payment system but also introduced the bad habits of overspending money. But, be careful when it starts allowing credit payments, as that might tempt people to spend more.”-[P13]
Our study uncovers clear trends in UPI adoption across user groups. Working professionals and business owners have seamlessly integrated UPI into their financial routines, using it for bill payments, invoicing clients, and managing business expenses. On the other hand, students from various age groups are adopting UPI for academic needs, such as paying tuition fees, buying books, and subscribing to online courses. Common spending trends, such as grocery shopping, cab bookings, and food delivery, were observed consistently across age groups and professions, demonstrating the broad utility of UPI for routine transactions and its transformative influence on spending behavior.
4.2.3. ATTITUDE TOWARDS FINANCIAL MANAGEMENT
Respondents held diverse perspectives on their financial discipline, with some expressing confidence in their ability to manage finances (rating 4 or 5) while others appeared less assured (giving ratings of 2 or 3). This shows that people have different levels of comfort with money matters.
Opinions about how UPI affects financial management were mixed. Some found it helpful, while others felt it complicated things. This suggests that the impact of UPI on personal finances varies depending on individual preferences and habits.
“Since I can refer to my balance and track my spending, controlling my expenses and budget has become easier. The same may not be the case for many of the population since the accessibility feature of UPI makes it harder for people to stick to a budget, which leads to overspending.”-[P14]
4.2.4. RECOMMENDATIONS FOR FINANCE TRACKING APPLICATIONS
Users expressed the need for visualization tools like graphical representations and pie charts to help them understand their spending patterns. The ability to track expenses over time and categorize them was also a common request. Users want insights into where their money is going, allowing them to make more informed financial decisions.
“I’d like my UPI app to have some specific features. It should allow me to restrict spending in certain categories, such as food, by using QR codes to track locations and deny transactions in those places. I want the app to generate a pie chart showing the average payments I make in different categories over a specific period, making tracking my expenses easier.”-[P15]
Responses from various professional backgrounds revealed that the specific needs of their professions often shape users’ expectations of financial management apps. Business professionals, including entrepreneurs, sought tools to manage business-related finances efficiently, such as tracking client payments and categorizing expenses. Lawyers expressed the need for streamlined record-keeping for legal practices, while doctors aimed to collect patient payments and related transactions effectively. Working-class professionals valued simplicity and ease of use in financial apps, emphasizing graphical representations and clear spending patterns. These insights underscore the importance of tailoring financial apps to meet the unique requirements of diverse professions, ensuring effective financial management across various occupational backgrounds.
“As a doctor, it would be beneficial if I could easily keep track of all the patients who have paid their fees. Checking transaction histories repeatedly can be quite a hassle. So, having an automatically generated list of patients who have made payments, with a clear categorization of whether the payment has been completed, would be incredibly convenient.”-[P16]
Some respondents went beyond traditional budgeting and desired financial education and guidance within the app. They emphasized the importance of learning to achieve financial stability and discipline through the application’s features.
“Besides handling expenses and transactions, it would be great if the app also offers guidance on achieving financial stability and discipline, perhaps by providing tips and educational content to help users make better financial decisions.”-[P17]
Respondents highlighted the importance of receiving notifications and alerts when approaching spending limits or exceeding budgeted amounts. This feature is a valuable tool to maintain financial discipline.
“An app where we can set limits for miscellaneous grocery shopping expenses and alert users about their financial condition. I’d appreciate having a daily spending limit feature that sends me reminders when I approach or reach my daily spending goals, helping me manage my finances more effectively.”-[P18]
5. Discussion
FEATURES
YEARLY SUBSCRIPTION COST
PROS
CONS
SUPPORTING PLATFORMS
FinArt (fin, 2016)
Automatically organizes transactional SMS into expenses, budget, and bill reminders, even when the app is closed.
Rs 799 (9.61 USD)
Data stored on Drive, Auto Tracking, spend analysis dashboards, and budget setting. Separates personal and business accounts for comprehensive financial management.
Setup complexity, ads persist after subscription, and occasional difficulty in SMS tracking.
ANDROID, PC, IOS
Money Manager (mm, 2020)
The app efficiently handles dual accounts, ensuring separate tracking of income and expenses. It offers offline/cloud backup and Data Export functionalities.
Free
Diverse payment method profiles, encompassing cash and cards, support competitive analysis. Easily create new profiles based on specific needs.
The application takes a long time if currencies are changed to optimize expenses and income. Most fees are non-editable.
ANDROID, IOS
Monefy (my, 2016)
The app is straightforward; the infographics make it vibrant and easy to use.
Rs 180 (2.17 USD)
Set bill reminders, manage multiple accounts and easily import data through Excel.
Frequent prompts for premium upgrades
ANDROID, PC ,IOS
Wallet (wal, 2014)
Users can use it for different currency rates worldwide. The Group Sharing option is also available in the application
Rs 700 (8.42 USD)
Monitor group expenses in the totals tab, track recent costs in the activity tab, and official API support available.
Accounts auto-disconnect; unable to modify the initial login currency.
ANDROID, IOS
Splitwise (spl, 2011)
App simplifies bill splitting , not for personal expense tracking. Form groups, manage shared bills, and track dues effortlessly.
Rs 499 (6 USD)
Clean UI. Manages multiple wallets, links bank accounts or Crypto wallets directly, and tracks transactions seamlessly (premium feature).
Tracking and analyzing personal finances can be challenging.
ANDROID, PC, IOS
Spendee (spe, 2013)
The app provides user-friendly expense and income tracking with easy-to-use tags for sorting costs efficiently.
Rs 900 (10.43 USD)
Effortlessly track expenses, share budgets, view spending analytics, and easily monitor debt progress.
Certain features are restricted behind a paywall.
ANDROID, PC, IOS
GoodBudget (gb, 2009)
Budget management application. The app lets you create and refill multiple budgets (envelopes) at intervals like monthly or weekly.
Rs 5817 (70 USD)
Create multiple wallets, set up recurring payments, track credit card and debt. Link bank accounts and enable group sharing (premium only).
Simple UI. Adding transactions is tedious, and there’s no API support.
ANDROID, PC, IOS
Table 1. Comparison of Personal Finance Tracking Apps
Answering RQ1 as mentioned in Section 1, the study results reveal a significant impact of UPI on Indian users’ spending habits. Most respondents (74.2%) accepted that UPI adoption led to increased spending, highlighting the UPI influence on user behavior. The data also indicates that 12% of respondents started using UPI after the demonetization in 2016 and , a 16% and 17.6% in 2020 and 2021 respectively, a period marked by significant economic and social changes. This suggests that external events, such as demonetization and pandemic-related lockdowns, may have been one of the causes of adoption of UPI, influencing spending patterns. Furthermore, 59.8% of respondents acknowledged exceeding their budget or overspending due to UPI usage. The qualitative analysis supports these findings. Overall, the evidence suggests that UPI has contributed to increased spending for most users. Users have increasingly turned to UPI for a wide range of transactions, including P2P (peer-to-peer) and P2M (peer-to-merchant) transactions, with P2P transactions taking the lead (met, 2023). With the integration of UPI as a payment method in multiple apps and places like street vendors, users are more likely to engage in impulse purchases, smaller transactions, and online shopping. Additionally, the qualitative analysis highlighted the psychological shift users experience, with the abstract nature of digital transactions (”just a number on the screen”) leading to impulsive spending. In some cases, regret and guilt hit people after they realize their spending by checking their bank balances or statements, compared to spending physical cash, highlighting the lack of tangibility associated with UPI payments. Answering RQ2 as mentioned in Section 1, user satisfaction levels with 91.5% of the survey respondents reporting satisfaction with the usage of UPI and 95.2% of the survey respondents found making payments via UPI convenient. The qualitative analysis underscores the convenience of UPI transactions, with users praising its speed, efficiency, and the elimination of the need for physical cash. The absence of processing fees, as opposed to credit and debit cards, is also highlighted as a positive factor contributing to user satisfaction. The research findings indicate that users perceive UPI as a secure and efficient payment method, eliminating concerns about carrying large sums of money. This positive perception of convenience and efficiency likely contributes to the widespread adoption and satisfaction with UPI, ultimately shaping users’ preferences for cashless transactions and influencing spending behavior. The research also notes that the cashless nature of UPI enhances business operations for professionals like lawyers, emphasizing the efficiency gained through expediting payments.
Answering RQ3 as mentioned in Section 1, the study identifies variations in UPI adoption across demographic groups, particularly in professions. Working professionals and business owners seamlessly integrate UPI into their financial routines, using it for various purposes such as bill payments and managing business expenses. On the other hand, students adopt UPI for academic needs like paying tuition fees. These differences highlight the importance of tailoring financial solutions to meet the specific needs of diverse user segments. Despite variations in adoption, common spending trends are observed across age groups and professions, demonstrating the broad utility of UPI. The findings suggest that the adoption of UPI is influenced by users’ professional backgrounds and financial needs, showcasing the platform’s adaptability to various user requirements.
5.1. RECOMMENDATION AND DESIRED FEATURE IN FINANCE TRACKING APPLICATION AND UPI
Our research insights highlight key recommendations for UPI platforms and financial app developers, aiming to elevate the digital payment experience. Users increasingly seek budgeting and expenditure limitation features within UPI apps, enabling them to set spending thresholds for specific categories and receive timely reminders to stay within financial goals. Visualization tools like graphical representations and pie charts should be integrated to enhance user comprehension, offering a clear overview of spending patterns over time.
In addition, there is a growing demand for educational content on responsible financial management embedded within the apps to empower users with financial literacy. Recognizing the diverse needs of different professions is crucial; for instance, business professionals may require tools to manage client payments and categorize expenses efficiently. Tailoring UPI solutions to cater to various user segments ensures accessibility and relevance.
5.2. FINANCIAL APP ANALYSIS IN INDIA
Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1. These applications offer diverse features tailored to meet individual financial requirements, ranging from automated organization of transactional data to robust expense-tracking capabilities. Users can leverage these tools to elevate their financial management skills. However, some notable gaps exist in meeting the recommendations outlined in Section 5.1. Specific features like detailed graphical representations of expenditure patterns and spending thresholds with reminders are unevenly available across platforms. Providing educational content on responsible financial management also varies, with certain apps lacking sufficient resources. Addressing these gaps requires finer control mechanisms for spending limits, real-time notifications, goal-based savings integration, AI-powered financial insights, customizable reporting, collaborative budgeting, integration with financial education platforms, and improvements in expense categorization. Enhancing user convenience should involve autonomously pulling expense data without manual input, ensuring a seamless experience while safeguarding user privacy. Balancing automation with privacy protection is crucial for advancing a comprehensive and empowering financial tracking solution for users.
In conclusion, while existing financial tracking apps offer valuable features, there is room for improvement in aligning with the recommendations outlined. Future developments should focus on bridging these gaps to provide users with a more holistic and empowering financial management experience.
Stakeholders, including app developers, government agencies, and users, stand to benefit from incorporating our suggestions into future financial tracking app developments. This comprehensive analysis addresses user needs and provides valuable insights for enhancing the overall financial management ecosystem, ultimately advancing effective and user-centric solutions.
5.3. FINANCIAL DISCIPLINE AND UPI USAGE
Our research shows the pivotal role of financial discipline in UPI usage. User interviews revealed a distinct contrast between those with strong financial discipline, harnessing UPI’s convenience without compromising financial stability, and those who occasionally succumbed to overspending. The latter group often cited a lack of financial literacy and self-control contributing to their spending behavior. The abstract nature of digital transactions in UPI can sometimes lead to impulsive purchases. Therefore, it is imperative to promote financial education and in helping the user understand their expenses and responsible spending practices alongside adopting digital payment systems like UPI. This two-pronged approach can empower users to maximize the advantages of UPI while safeguarding their financial well-being.
5.4. FUTURE IMPLICATIONS AND POLICY CONSIDERATION
The research outcomes hold significant implications for various stakeholders in India. Policymakers can use this information to design regulations that encourage responsible UPI usage and protect users from potential scams. Businesses and financial institutions should adapt their services to accommodate the diverse needs of users from various age groups and professions, enhancing their overall user experience. Additionally, educational initiatives can focus on improving financial literacy and promoting responsible spending habits among UPI users. This research emphasizes the transformative power of UPI in reshaping spending behavior, making it crucial for all involved parties to adapt and harness its full potential for the benefit of Indian consumers and the broader economy. Many other countries, like France, Singapore, etc., have started embracing UPI to promote digital payments (glo, 2023). With the recent developments by NPCI like UPI integration with Rupay Credit Card (rup, 2012), tap and pay feature with Near Field Communication (NFC) (nfc, 2004), UPI Lite (upi, 2022) and UPI LiteX (upi, 2023) for offline payments, the infrastructure of UPI is being aimed to scale to 100 billion monthly transactions, making it a greater force of revolution in the Indian financial landscape.
6. Conclusion
In this paper, we conducted a user study to understand the influence of UPI on spending behavior among Indian users. We studied how UPI has enabled convenience and seamless financial transactions, overshadowing cash transactions. UPI has played a pivotal and remarkable role in India’s push for a cashless economy and is also benefiting India’s GDP. Our study culminates by aligning findings with research questions and user recommendations, providing actionable insights for enhancing UPI app features and facilitating effective expenditure management. This research has broader implications for policymakers, businesses, and financial institutions, highlighting the need to adapt and promote responsible UPI usage to maximize its transformative potential in the Indian economic landscape.